OpenStudy (pphalke):

8x-(5x+5)=4 solve will medal

OpenStudy (anonymous):

Distribute the negative. Then collect like terms.

OpenStudy (aveline):

If it helps, you might want to rewrite the equation: 8x-1(5x+5)=4 Distributing a negative is technically multiplying by 1 So, 8x-(1*5x)-(5)=4 8x-5x-5=4 3x-5=4 3x=9 x=3

OpenStudy (moji7029):

I would go with Zelda